# Environmentally Sustainable Software Design and Development: A Systematic Literature Review

### v1

@@ 3.2. Reference models
Raisian et al. (Raisian et al., 2021) propose green and sustainable software measurements as a reference model for software developers that aim at targeting sustainability by considering the environmental dimension (i.e., energy and carbon footprints in processing and storage), the economic dimension (i.e., usability, available time and costs), and the social dimension (i.e., support tools, human well-being, performance). To achieve the above, they aim at identifying suitable hierarchical measurements that are to be evaluated, documented, and assessed for continuous software enhancement. Still with a hierarchical approach to continuous software improvement, Penzenstadler and Femmer (Penzenstadler and Femmer, 2013) provide a reference model definition based on different levels of granularity, for both software architects and sustainability engineers. After defining sustainability metrics and their targets, the proposed model foresees matching software life cycle activities to their (positive/negative) impacts on energy or carbon emissions, and identifying responsible personnel for each activity.
@@
Becker et al. (Becker et al., 2016) propose a reference model that allows product owners, software providers, developers and sustainability engineers to compare two equivalent software systems, designed and implemented with and without considering sustainability, respectively. The model aligns sustainable goals with those of the system, determines project scope and system boundaries for any stakeholder, makes the risk analysis for each sustainability dimension, and finally, implements requirements and specifications based on templates including checklists for sustainability criteria and standards compliance in all the interrelated dimensions. Lago and Jansen (Lago and Jansen, 2010) propose a reference model that supports the iterative and incremental way of operating for software developers and end-users. Such model is used to drive the implementation of two Web services, i.e., the Environmental Strategies as-a-Service and the Green Metrics as-a-Service, useful to make the environmental impact of software services visible by continuously monitoring it, giving feedback and tailoring software based on the available data to reduce energy and carbon footprints.
@@
We conclude by presenting the GREENSOFT reference model, initially defined by Naumann et al. (Naumann et al., 2011), which led to many other research efforts. Following a holistic approach, GREENSOFT is intended for all stakeholders, i.e.,software architects, developers, IT operators, and end-users. The model seconds the Agile software development model and considers first (green-IT), second (green-by-IT) and third order (systemic) effects related to designing, implementing, using and decommissioning software products. It foresees the definition of ad hoc sustainability criteria and metrics, the analysis of procedure models involved in the development, administration, purchase and use of software, and the adoption of recommendations and tools targeting all stakeholders. Agile sprints are enhanced by including sustainability reviews and retrospectives that focus on measured metrics and criteria, monitored via a sustainability journal throughout requirements elicitation, design, implementation and testing. These steps are mentioned by Dick et al. (Dick et al., 2013) when considering the sustainable Agile development stage.
@@
Retaking the process assessment model of GREENSOFT, Johann et al. (Johann et al., 2011) propose a reference model to develop sustainable software products, while mitigating negative impacts and amplifying positive impacts based on a knowledge base shared among different stakeholders. Their proposal includes a first implementation of recommendations and tools in the form of checklists and suggestions. Also Dick et al. (Dick et al., 2010) extends GREENSOFT by relying on the definition of checklists and best practices realised through interviewing stakeholders, with a focus on software development and project management. Particularly, they highlight how sustainability decisions taken at a set step of the software life cycle only show their impact during the next one.
@@
Later on, Mahmoud and Ahmad (Mahmoud and Ahmad, 2013) re-consider the GREENSOFT model and discuss how the software itself can be used as a tool to aid and promote fine-grained green computing for operating systems, frameworks, performance monitoring, programs written for energy management purposes, and virtualisation. Their mode l also accounts for fault-tolerance and failure management. Finally, following GREENSOFT and the ISO/IEC 12207 standard,444Systems and software engineering — Software lifecycle processes, International Organization for Standardization, 2008 Naumann et al. (Naumann et al., 2015) define criteria to identify software quality aspects as per a process-centric model, inherent both to the final product and the implemented process, through life cycle assessment. Their model addresses software developers, sustainability engineers and end-users and is given in a full-fledged Agile version following the whole life cycle as well as in a lightweight version focussing on process assessment and sustainability reviews and previews.
@@ 3.3. Measurement Solutions
This section focuses on the portion of the analysed corpus about measurement solutions for software sustainability, which take the form of estimate techniques and assistive tooling. Their aim is to measure or estimate energy consumption and carbon emissions related to software design and development so as to make informed choice throughout the software life cycle.
@@
Energy-awareness. Starting from those approaches that only focus on energy consumption, Noureddine et al. (Noureddine et al., 2012) develop a real-time software energy monitoring framework intended for both software developers and end-users. Such solution reports estimates on energy consumption by relying on a service-oriented architecture made of a monitoring sensor, a database, and a life-cycle monitoring module. Specifically, the individual components are isolated and managed using start/stop options, thus providing flexibility while continuously monitoring the deployment hardware (i.e., processors, network interfaces). The approach is illustrated over different programs written in different programming languages.
@@
Focussing instead on the HPC domain, Kipp et al. (Kipp et al., 2011) introduce an estimate technique intedend for sustainability engineers. They propose Green Performance Indicators as a supplement to Key Performance Indicators encompassing factors such as IT virtual/real resource usage, service lifecycle, approximate environmental impact, and organisational considerations. They consequently assess the sustainability across various levels of system architecture - from the abstract view to the business logic and up to the entire IT service centre. The methodology involves monitoring clusters of metrics for energy-intensive applications by first analysing single jobs and observing the relation between energy consumption and execution time for each run. Such data is then input to a constraint-based formulation to optimise resource usage and energy consumption.
@@
Carbon-awareness. Sharma et al. (Mehra et al., 2022) devise an assistive tooling for software developers to assess the energy efficiency of their projects throughout the development life cycle. This tool provides a standard set of metrics to evaluate the energy consumption impact of different design decisions at each stage of the project. The approach is based on a series of questions related to project characteristics, e.g., programming languages, used third-party libraries, cloud providers, technologies, processes, and teams’ organisation. Ultimately, the tool provides an overall sustainability score and recommendations to further improve software sustainability. To also promote responsibility, awareness and transparency, Kern et al. (Kern et al., 2015) develop a model- and tool-based calculation method intended for software developers and product owners to measure the carbon footprint of their products according to international standards. The strategy involves collecting data throughout the software development life cycle to continuously monitor carbon emissions.
@@
Energy-awareness and carbon-awareness. This last section contains a large number of surveyed studies, scattered across a plethora of different domains. Starting from generic software, Pa et al. (Pa et al., 2017) propose an assistive tool to help developers in understanding and implementing sustainable software by measuring software energy and carbon consumption. The prototype feautes online accessible web-based dashboards to visualise sustainability reports and suggestions for defining best trade-off solutions between database organisation, network usage and human resources. The tool can also track historical data to monitor project evolution over time. Simon et al. (Simon et al., 2023) provide a comprehensive estimate technique based on graph modelling regarding the environmental impact of software life cycle for project owners. They have nodes representing life cycle steps and edges describing different environmental impacts based on direct observations. Life cycle assessment follows a lightweight bottom-up approach to identify stakeholders’ responsibility and actionable insights and to identify hot spots for potential improvement. Still targeting project owners, Moshnyaga (Moshnyaga, 2013a, b) analyses the total energy consumption associated with the software throughout its entire development lifecycle and evaluates its contribution to carbon emissions, also accounting for hardware-specific emissions. The proposed methodology also relies on rapid prototyping to build software following an iterative three-step procedure consisting of business model definition, analysis (revision) of risk, and development plan (update). The same authors also tailored their proposal specifically targeting mobile applications (Moshnyaga, 2013c), along with both project owners and software developers.
@@
Speaking of the ML domain, based on Mehra et al. (Mehra et al., 2022), Sikand et al. (Sikand et al., 2023) present a standard metric set for measuring the energy efficiency of artificial intelligence systems throughout the development lifecycle. They also call attention to harmful aspects associated with red AI, e.g., the exponential growth in data, model complexity, large models, increased number of experiments, and the relentless pursuit of higher performance. Extending their previous effort (Kipp et al., 2011), Kipp et al. (Kipp et al., 2012) focus instead on HPC by introducing Green Performance Indicators to complement the traditional KPIs, by enabling a comprehensive evaluation of the environmental impact and sustainability classification of IT service centres by examining various hierarchical system levels, from the application level to the middle-ware and up to the infrastructure.
@@ 3.4. Software Improvement Techniques
Energy-awareness. Kocak and Alptekin (Koçak and Alptekin, 2019) propose a multi-criteria decision-making technique for generic software, as part of the analysis and design of environmentally sustainable software. Their technique aims at balancing between the needs of end-users, who want energy-efficient software at a low cost, and software providers, who offer products at a maximum quality level. The proposal addresses the conflicting interests by determining demand and supply functions for decision-making entities that are considered to be rational, intelligent, and having a comprehensive plan of action for any situation that may arise. The proposal also discusses implementing a strategic interaction based on the inter-dependencies of software product properties and attributes. They also align their proposal to the ISO/IEC 25010 standards, including attributes such as functional suitability, reliability, performance efficiency, and usability. The authors then propose a method to translate these requirements into terms of price, quality, and utility.
@@
Focussing on mobile applications, Le Goaër (Goaër, 2021) provides a code scanning tool for the Android Studio IDE. The tool performs white-box static analyses to identify green code smells and associated potential energy optimisation. The scanning tool is designed for software developers who use the Android Lint framework. Although individual energy savings may be small for single Android devices, the cumulative impact across billions of devices is significant. Different green bugs are matched to their severity, priority, and category, along with candidate single-click fixes (e.g., rely on dark UI, data compression, timeout wave locks). Still in the mobile landscapes, Beghoura et al. (Beghoura et al., 2014) propose instead a black box analysis for software developers that rely on a supervised vector machine model to create specific energy consumption models for different types of devices. They investigate the energy consumption patterns of hardware when running specific applications, accounting for various factors such as hardware configuration, product design, and battery ageing. More specifically, the study aims at understanding non-linear energy consumption components, such as changes in central processing unit power consumption based on frequency and voltage.
@@
Moving to the field of blockchains, Lasla et al. (Lasla et al., 2022) introduce a distributed protocol, which defines a green proof-of-work algorithm. The implementation comprises two rounds. The first involves all network nodes in the mining process, while the second involves a smaller set called runners-up and reduces energy consumption. The energy spent to mine one block is used to select a small number of miners to exclusively mine the next block. The protocol also considers and balances among blockchain fork frequency, security and decentralisation level.
@@
Carbon-awareness. Only the proposal by Karunakaran and Rao (Karunakaran and Rao, 2013), intended for project owners, provides a model-driven technique that solely consider the carbon footprint of software. They propose to exploit coloured Petri Nets that embed time constraints and hierarchical implementation to devise and configure optimal policies of a waterfall software development life cycle. Their model accounts for machine, software and resource levels to simulate programming, testing and debugging of software projects and estimate the associated carbon emissions.
@@
Energy-awareness and carbon-awareness. Hoesch-Klohe et al. (Hoesch-Klohe et al., 2010) provide a model-driven technique which leverages business process models for their capability to capture the relations between resources and activities, specifically focusing on carbon emission impact, through textual assertions. Their technique involves identifying all process designs, ordering them based on their abstraction level, and then systematically determining a suitable association to computing resources accounting both for carbon emissions and energy consumption. Lago (Lago, 2019) devises a decision-making technique intended for software architects, based on architectural designs and assessments for sustainability. In the specific, decision maps capture the impacts of sustainability dimensions as immediate, enabling, and systemic, and categorise their effects as positive, negative, or undecided. Decision maps are used to explore the software design space by addressing multiple sustainability perspectives at the same time; they can be used to drive the decision-making process related to software design, while refining it iteratively.
@@
Focussing on cyber-physical systems, Heithoff et al. (Heithoff et al., 2023) present instead a model-based approach for project owners and software developers. By relying on digital twins and digital shadow systems, they monitor the sustainability of the actual real-world software. Digital twins are designed to be smaller in size than the original system and allow for faster analyses through simulated behaviour. Their prototype also enables monitoring energy consumption peaks in comparison to current workloads and predefined levels. With cloud computing developers as their target, Jebraeil et al. (Jebraeil et al., 2017) propose a model-driven technique called g-UML. They extend UML diagrams with contextual information about data centre and network operations, i.e., power generation, cooling and servers consumption, data communication footprint. This enables the efficient identification of energy-related issues and the organisation of power-related components into an architectural view. Resulting gUML diagrams can assist software developers in writing energy-efficient software, enabling collaboration between hardware and software architects. Addressing software developers, Fakhar et al. (Fakhar et al., 2012) combine a white-box static analysis technique for energy-efficient software compilation with a black-box analysis technique for effective task scheduling. The compiler translates source code in an energy-efficient format, optimises binary code through green and sustainable strategies, and provides hints for energy-efficient coding. On the other hand, the scheduler utilizes a greedy algorithm to efficiently make use of cloud resources, selecting and allocating machines based on user requirements while aiming at using as many processing cores in a machine as possible.
@@
Moving from cloud to edge computing settings, Zhang et al. (Zhang et al., 2023) present an optimisation technique tailored for software providers operating in a network that offers smart services. They propose the use of an evaluation framework that combines dynamic energy trading with task allocation optimisation for federated edge intelligence networks so as to reduce carbon emissions throughout the software life cycle. At design time, they mainly focus on computing carbon emissions related to data-generating devices. At development time, they account for emissions due to pre-training, testing, parameter selection and optimisation, based on specified service requirements and usage scenarios. At deployment time they address model distribution, while the disposal phase focuses on generalisation, modification, and transfer issues.
@@
In line with Le Goaër (Goaër, 2021), Le Goaër et al. (Goaër, 2023; Goaër and Hertout, 2022) presents a white-box static computer-aided customisable code inspection tool for Android software developers, named ecoCode. This tool is designed to identify and refactor code structures that are inefficient in terms of energy consumption (thus impacting on carbon emissions). By emphasising energy efficiency as a quality attribute, ecoCode aims at promoting the concept of green and sustainable code, akin to the way security and maintainability are enforced in clean code. Particularly, ecoCode performs a scan of project repositories identify environmental issues, which are then displayed on a web dashboard.
@@
Finally, in the domain of blockchains, Alofi et al. (Alofi et al., 2023, 2022) propose a search-based software engineering technique for addressing energy and carbon consumption concerns. The proposal is intended for software architects and sustainability engineers and employs evolutionary meta-heuristics to identify subset of miners that maximise trust levels in the system, maintain decentralisation and reduce energy consumption and carbon emissions. They propose to implement the system as an autonomic MAPE-K loop that ensures enactment of taken decisions.

### v2

@@ 3.2. Reference models
Raisian et al. (Raisian et al., 2021) propose green and sustainable software measurements as a reference model for software developers that aim at targeting sustainability by considering the environmental dimension (i.e., energy and carbon footprint in processing and storage), the economic dimension (i.e., usability, available time and costs), and the social dimension (i.e., support tools, human well-being, performance). To achieve the above, they aim at identifying suitable hierarchical measurements that are to be evaluated, documented, and assessed for continuous software enhancement. Still with a hierarchical approach to continuous software improvement, Penzenstadler and Femmer (Penzenstadler and Femmer, 2013) provide a reference model definition based on different levels of granularity, for both software architects and sustainability engineers. After defining sustainability metrics and their targets, the proposed model foresees matching software lifecycle activities to their (positive/negative) impacts on energy or carbon emissions, and identifying responsible personnel for each activity.
@@
Becker et al. (Becker et al., 2016) propose a reference model that allows product owners, software providers, developers and sustainability engineers to compare two equivalent software systems designed and implemented with and without considering sustainability, respectively. The model aligns the sustainability goals with those of the system, defines the project scope and system boundaries for any stakeholder, makes the risk analysis for each sustainability dimension, and finally, implements requirements and specifications based on templates including checklists for sustainability criteria and standards compliance in all the interrelated dimensions. Lago and Jansen (Lago and Jansen, 2010) propose a reference model that supports the iterative and incremental way of operating for software developers and end-users. The model is used to drive the implementation of two Web services, i.e., the Environmental Strategies as-a-Service and the Green Metrics as-a-Service, useful to make the environmental impact of software services visible by continuously monitoring it, giving feedback and tailoring software based on the available data to reduce energy and carbon footprints.
@@
We conclude by presenting the GREENSOFT reference model, originally defined by Naumann et al. (Naumann et al., 2011), which led to many other research efforts. Following a holistic approach, GREENSOFT is intended for all stakeholders, i.e.,software architects, developers, IT operators, and end-users. The model seconds the Agile software development model and considers first (green-IT), second (green-by-IT) and third order (systemic) effects related to designing, implementing, using and decommissioning software products. It foresees the definition of ad hoc sustainability criteria and metrics, the analysis of process models for the development, administration, purchase and use of software, and the adoption of recommendations and tools for all stakeholders. Agile sprints are enhanced by including sustainability reviews and retrospectives that focus on measured metrics and criteria, monitored via a sustainability journal throughout requirements elicitation, design, implementation and testing. These steps are mentioned by Dick et al. (Dick et al., 2013) when considering the sustainable Agile development stage.
@@
Retaking the process assessment model of GREENSOFT, Johann et al. (Johann et al., 2011) propose a reference model to develop sustainable software products, while mitigating negative impacts and amplifying positive impacts based on a knowledge base shared among different stakeholders. Their proposal includes a first implementation of recommendations and tools in the form of checklists and suggestions. Dick et al. (Dick et al., 2010) also extend GREENSOFT by relying on the definition of checklists and best practices realised through stakeholder interviews, with a focus on software development and project management. Particularly, they highlight how sustainability decisions taken at a set step of the software lifecycle only show their impact during the next one.
@@
Later, Mahmoud and Ahmad (Mahmoud and Ahmad, 2013) re-consider the GREENSOFT model and discuss how software itself can be used as a tool to aid and promote fine-grained green computing for operating systems, frameworks, performance monitoring, programs written for energy management purposes, and virtualisation. Their model also accounts for fault-tolerance and failure management. Finally, following GREENSOFT and the ISO/IEC 12207 standard,444Systems and software engineering — Software lifecycle processes, International Organization for Standardization, 2008 Naumann et al. (Naumann et al., 2015) define criteria to identify software quality aspects as per a process-centric model, inherent both to the final product and the implemented process, through lifecycle assessment. Their model addresses software developers, sustainability engineers and end-users and is given in a full-fledged Agile version following the whole lifecycle as well as in a lightweight version focussing on process assessment and sustainability reviews and previews.
@@ 3.3. Measurement Solutions
This section focuses on the portion of the analysed corpus about measurement solutions for software sustainability, which take the form of estimate techniques and assistive tooling. Their aim is to measure or estimate carbon emissions related to software design and development so as to make informed choice throughout the software lifecycle.
@@
Indirect. Starting from those approaches that indirectly consider carbon emissions, Noureddine et al. (Noureddine et al., 2012) develop a real-time software energy monitoring framework intended for both software developers and end-users. Such solution reports estimates on energy consumption by relying on a service-oriented architecture consisting of a monitoring sensor, a database, and a lifecycle monitoring module. Specifically, the individual components are isolated and managed using start/stop options, thus providing flexibility while continuously monitoring the deployment hardware (i.e., processors, network interfaces). The approach is illustrated over different programs written in different programming languages.
@@
Focussing instead on the HPC domain, Kipp et al. (Kipp et al., 2011) introduce an estimate technique intended for sustainability engineers. They propose Green Performance Indicators as a supplement to Key Performance Indicators encompassing factors such as the use of virtual/real IT resources, the service lifecycle, approximate environmental impact, and organisational considerations. They consequently assess sustainability at different levels of a system architecture - from the abstract view to the business logic and up to the entire IT service centre. The methodology involves monitoring clusters of metrics for energy-intensive applications by first analysing single jobs and observing the relation between energy consumption and execution time for each run. Such data is then input to a constraint-based formulation to optimise resource usage and energy consumption.
@@
Direct. Sharma et al. (Mehra et al., 2022) devise an assistive tooling for software developers to assess the carbon efficiency of their projects throughout the development lifecycle. This tool provides a standard set of metrics to evaluate the energy consumption impact of different design decisions at each stage of the project. The approach is based on a series of questions related to project characteristics, e.g., programming languages, used third-party libraries, cloud providers, technologies, processes, and teams’ organisation. Ultimately, the tool provides an overall sustainability score and recommendations to further improve software sustainability. To also promote responsibility, awareness and transparency, Kern et al. (Kern et al., 2015) develop a model- and tool-based calculation method intended for software developers and product owners to measure the carbon footprint of their products according to international standards. The strategy involves collecting data throughout the software development lifecycle to continuously monitor carbon emissions.
@@
Hybrid. This last section contains a large number of surveyed studies, scattered across a plethora of different domains. Starting from generic software, Pa et al. (Pa et al., 2017) propose an assistive tool to help developers in understanding and implementing sustainable software by estimating its carbon footprint, also based on its energy consumption. The prototype features online accessible web-based dashboards to visualise sustainability reports and suggestions for defining best trade-off solutions between database organisation, network usage and human resources. The tool can also track historical data to monitor project evolution over time. Simon et al. (Simon et al., 2023) provide a comprehensive estimate technique based on graph modelling regarding the environmental impact of software lifecycle for project owners. They have nodes representing lifecycle stages and edges describing different environmental impacts based on direct observations. Lifecycle assessment follows a lightweight bottom-up approach to identify stakeholders’ responsibility and actionable insights and to identify hot spots for potential improvement. Still targeting project owners, Moshnyaga (Moshnyaga, 2013a, b) analyses the total energy consumption associated with the software throughout its entire development lifecycle and evaluates its contribution to carbon emissions, also accounting for hardware-specific emissions. The proposed methodology also relies on rapid prototyping to build software following an iterative three-step process consisting of business model definition, risk analysis (revision), and development plan (update). The same authors also tailored their proposal specifically targeting mobile applications (Moshnyaga, 2013c), along with both project owners and software developers.
@@
Speaking of the ML domain, based on Mehra et al. (Mehra et al., 2022), Sikand et al. (Sikand et al., 2023) present a standard set of metrics for measuring the energy efficiency of artificial intelligence systems throughout the development lifecycle. They also call attention to harmful aspects associated with red AI, e.g., the exponential growth in data, model complexity, large models, increased number of experiments, and the relentless pursuit of higher performance. Extending their previous effort (Kipp et al., 2011), Kipp et al. (Kipp et al., 2012) instead focus on HPC by introducing Green Performance Indicators to complement the traditional KPIs, by enabling a comprehensive evaluation of the environmental impact and sustainability classification of IT service centres by examining various hierarchical system levels, from the application level to the middle-ware and up to the infrastructure.
@@ 3.4. Software Improvement Techniques
Indirect. Kocak and Alptekin (Koçak and Alptekin, 2019) propose a multi-criteria decision-making technique for generic software, as part of the analysis and design of carbon efficient software. Their technique aims at balancing between the needs of end-users, who want energy-efficient software at a low cost, and software providers, who offer products at a maximum quality level. The proposal addresses the conflicting interests by determining demand and supply functions for decision-making entities that are considered rational, intelligent, and having a comprehensive plan of action for any situation that may arise. The proposal also discusses the implementation of a strategic interaction based on the interdependencies of software product properties and attributes. They also align their proposal with the ISO/IEC 25010 standards, including attributes such as functional suitability, reliability, performance efficiency, and usability. The authors then propose a method to translate these requirements into terms of price, quality, and utility.
@@
Focussing on mobile applications, Le Goaër (Goaër, 2021) provides a code scanning tool for the Android Studio IDE. The tool performs white-box static analyses to identify green code smells and associated potential energy optimisation. The scanning tool is designed for software developers who use the Android Lint framework. Although individual energy savings may be small for single Android devices, the cumulative impact across billions of devices is significant. Different green bugs are matched to their severity, priority, and category, along with candidate single-click fixes (e.g., rely on dark UI, data compression, timeout wave locks). Still in the mobile landscape, Beghoura et al. (Beghoura et al., 2014) instead propose a black box analysis for software developers that relies on a supervised vector machine model to create specific energy consumption models for different types of devices. They investigate the energy consumption patterns of hardware when running specific applications, accounting for various factors such as hardware configuration, product design, and battery ageing. More specifically, the study aims at understanding non-linear energy consumption components, such as changes in CPU power consumption as a function of frequency and voltage.
@@
Moving to the field of blockchains, Lasla et al. (Lasla et al., 2022) introduce a distributed protocol, that defines a green proof-of-work algorithm. The implementation comprises two rounds. The first involves all network nodes in the mining process, while the second involves a smaller set of nodes, called runners-up. The energy consumption is reduced since the same energy spent to mine one block is also used to select a small number of miners to exclusively mine the next block. The protocol also considers and balances among usual blockchain tradeoffs such as fork frequency, security and decentralisation level.
@@
Direct. Only the proposal by Karunakaran and Rao (Karunakaran and Rao, 2013), intended for project owners, provides a model-driven technique that solely considers the carbon footprint of software. They propose to exploit coloured Petri Nets that embed time constraints and hierarchical implementation to devise and configure optimal policies of a waterfall software development lifecycle. Their model accounts for machine, software and resource levels to simulate programming, testing and debugging of software projects and estimate the associated carbon emissions.
@@
Hybrid. Hoesch-Klohe et al. (Hoesch-Klohe et al., 2010) provide a model-driven technique which leverages business process models for their capability to capture the relations between resources and activities, specifically focusing on carbon emission impact, through textual assertions. Their technique involves identifying all process designs, ordering them based on their abstraction level, and then systematically determining a suitable association to computing resources accounting both for carbon emissions and energy consumption. Lago (Lago, 2019) devises a decision-making technique intended for software architects, based on architectural designs and assessments for sustainability. In the specific, decision maps capture the impacts of sustainability dimensions as immediate, enabling, and systemic, and categorise their effects as positive, negative, or undecided. Decision maps are used to explore the software design space by addressing multiple sustainability perspectives at the same time; they can be used to drive the software design decision process, while refining it iteratively.
@@
Focussing on cyber-physical systems, Heithoff et al. (Heithoff et al., 2023) instead present a model-based approach for project owners and software developers. By relying on digital twins and digital shadow systems, they monitor the sustainability of the actual real-world software. Digital twins are designed to be smaller in size than the original system and allow for faster analyses through simulated behaviour. Their prototype also enables monitoring energy consumption peaks in comparison to current workloads and predefined levels. With cloud computing developers as their target, Jebraeil et al. (Jebraeil et al., 2017) propose a model-driven technique called g-UML. They extend UML diagrams with contextual information about data centre and network operations, i.e., power generation, cooling and servers consumption, data communication footprint. This enables the efficient identification of energy-related issues and the organisation of power-related components into an architectural view. Resulting gUML diagrams can assist software developers in writing energy-efficient software, enabling collaboration between hardware and software architects. Addressing software developers, Fakhar et al. (Fakhar et al., 2012) combine a white-box static analysis technique for energy-efficient software compilation with a black-box analysis technique for effective task scheduling. The compiler translates source code in an energy-efficient format, optimises binary code through green and sustainable strategies, and provides hints for energy-efficient coding. On the other hand, the scheduler utilizes a greedy algorithm to efficiently make use of cloud resources, selecting and allocating machines based on user requirements while aiming at using as many processing cores in a machine as possible.
@@
Moving from cloud to edge computing settings, Zhang et al. (Zhang et al., 2023) present an optimisation technique tailored for software providers operating in a network that offers smart services. They propose the use of an evaluation framework that combines dynamic energy trading with task allocation optimisation for federated edge intelligence networks to reduce carbon emissions throughout the software lifecycle. At design time, they mainly focus on computing carbon emissions related to data-generating devices. At development time, they account for emissions due to pre-training, testing, parameter selection and optimisation, based on specified service requirements and usage scenarios. At deployment time, they address model distribution, while, at disposal time, they focus on generalisation, modification, and transfer issues.
@@
In line with Le Goaër (Goaër, 2021), Le Goaër et al. (Goaër, 2023; Goaër and Hertout, 2022) present a white-box static, computer-aided, customisable code inspection tool for Android software developers, named ecoCode. This tool is designed to identify and refactor code structures that are inefficient in terms of energy consumption (thus impacting on carbon emissions). By emphasising energy efficiency as a quality attribute, ecoCode aims at promoting the concept of green and sustainable code, akin to the way security and maintainability are enforced in clean code. Particularly, ecoCode performs a scan of project repositories to identify environmental issues, which are then displayed on a web dashboard.
@@
Finally, in the domain of blockchains, Alofi et al. (Alofi et al., 2023, 2022) propose a search-based software engineering technique for addressing energy and carbon consumption concerns. The proposal is intended for software architects and sustainability engineers and employs evolutionary meta-heuristics to identify subsets of miners that maximise trust levels in the system, maintain decentralisation and reduce energy consumption and carbon emissions. They propose to implement the system as an autonomic MAPE-K loop that ensures enactment of taken decisions.